OpenStudy (cookiemonster18):

1.Bradley and Ryan are having a contest to see who can collect the most bumper stickers. Bradley started with 23 bumper stickers and plans to add 4 each month. Ryan has 9 bumper stickers and plans to add 6 each month. After how many months will they have the same number of bumper stickers? A.10 months B.2 months C.7 months D.51 months

OpenStudy (cookiemonster18):

I think it's C.7 months am I correct @FortyTheRapper

OpenStudy (kendricklamar2014):

C. 7 is correct

OpenStudy (kendricklamar2014):

Here is how I did it: So you have 2 equations Bradley \[23 + 4m\] Ryan \[9 + 6m\] M is Months and you want to figure out when these 2 will be equal...so set the 2 equations equal to each other. \[23 + 4m = 9 + 6m\] \[m = 7\]